# Fine-Grained Domain Generalization with Feature Structuralization

### v1

@@ Abstract
Fine-grained domain generalization (FGDG) is a more challenging task than traditional DG tasks due to its small inter-class variations and relatively large intra-class disparities. When domain distribution changes, the vulnerability of subtle features leads to a severe deterioration in model performance. Nevertheless, humans inherently demonstrate the capacity for generalizing to fine-gout-of-distribution data, leveraging structured multi-granularity knowledge that emerges from discerning both the commonality and specificity within categories. Likewise, we propose a Feature Structuralized Domain Generalization (FSDG) model, wherein features experience structuralization into common, specific, and confounding segments, harmoniously aligned with their relevant semantic concepts, to elevate performance in FGDG. Specifically, feature structuralization (FS) is accomplished through joint optimization of five constraints: a decorrelation function applied to disentangled segments, three constraints ensuring common feature consistency and specific feature distinctiveness, and a prediction calibration term. By imposing these stipulations, FSDG is prompted to disentangle and align features based on multi-granularity knowledge, facilitating robust subtle distinctions among categories. Extensive experimentation on three benchmarks consistently validates the superiority of FSDG over state-of-the-art counterparts, with an average improvement of 6.2% in terms of FGDG performance. Beyond that, the explainability analysis on explicit concept matching intensity between the shared concepts among categories and the model channels, along with experiments on various mainstream model architectures, confirms the validity of FS.
@@ I Introduction
In FGDG tasks, fine-grained categories exhibit multi-granularity structures, built upon the commonalities and specificities between categories. This aspect is overlooked by previous DG methods. In contrast, humans exhibit a higher level of generalizability by leveraging these structured semantics for learning and recognizing. According to family resemblances [9], feature integration theory [10], and lexical structure [11] in cognitive psychology, when learning to recognize objects, people focus primarily on the salient targets and inductively disentangle the commonalities and specificities among them. Commonalities represent basic visual concepts, whereas specificities correspond to individual conceptual extensions. Both serve as the foundational elements for object classification and vary across distinct granularity levels [12, 13, 14]. For instance, as Fig. 1 shows, four animals can be reclassified into categories of Equidae and Bovidae based on their commonalities and specificities from a higher $c_{1}$ granularity level, respectively. At $c_{1}$ level, zebra and horse share certain characteristics, such as long necks, while manifesting specificities at fine-grained level $f_{0}$. It can be observed that the common and specific features, which constitute the intrinsic characteristics of objects, can be facilitated by incorporating multi-granularity knowledge. Understanding the commonalities and specificities between categories enhances the generalizability, as both focus more on the intrinsic features of the objects themselves.
@@
However, a key challenge arises: how to explicitly embed these structured commonalities and specificities into DL models. In particular, it is challenging to determine what semantic basis can be used for constructing clusters of commonalities and specificities, and how to constrain them within their respective semantic clusters as literature has discussed the difficulty of learning decoupling semantics in the absence of sufficient additional knowledge constraints [15].
@@
To address this challenge, we propose a Feature Structuralization (FS) framework employing multi-granularity knowledge as an additional constraint for semantic disentanglement and alignment. Researches in the explainability field show that features or tokens have their own semantic meanings [16, 17, 18, 19], which implies that commonalities and specificities can be disentangled from the learned features. Besides, acknowledging that recognition tasks predominantly seek to discern target objects in an image that encompasses target semantics and confounding, we disentangle the total semantical features learned from an image into three components, namely commonality, specificity, and confounding. Furthermore, the representations of commonalities and specificities are constrained within their respective semantic clusters with the aid of a multi-granularity knowledge structure. The collaboration between both facilitates the network in learning category invariance while preserving learning discrimination. Consequently, FS optimizes the fine-grained invariant representation capabilities as both focus more on the attributes of the objects themselves rather than on the other spurious correlations.
@@
To formulate FS, we devise a pipeline consisting of two essential steps: disentanglement and alignment. We disentangle the three parts according to their channel indices. A decorrelation optimization function is employed to minimize the mutual information between them. Three similarity-based losses are then proposed for the alignment of common and specific parts with the aid of multi-granularity knowledge. Two of these are leveraged to pull commonalities closer across diverse granularities, while the third serves to create distance among the specific components. Furthermore, a Feature Structuralized Domain Generalization model (FSDG) is constructed by incorporating a prediction calibration method. Three model variations are expanded in conjunction with Convolutional Neural Network (CNN), Transformer, and Multi-Layer Perceptron (MLP). The contributions of this work are as follows:
@@
(1) We propose an FS framework to address a more difficult FGDG challenge, drawing inspiration from cognitive psychology. This novel approach integrates data and knowledge to structure the learned features into common, specific, and confounding parts according to their channel indexes, which is in line with the analysis of explainability.
@@
(2) We construct a feature-structuralized model FSDG by formulating FS into a disentanglement and alignment pipeline. A disentanglement enhancement function is derived to reduce the correlation between the three segments. Three similarity-based constraints integrating multi-granularity knowledge are then designed to achieve semantic alignment for the common and specific components. To the best of our knowledge, FSDG is a pioneer tailored to address FGDG issues.
@@
(3) The refined CRP method acquires the most relevant concepts for every category, showing that FS significantly increases the explicit concept matching intensity between the shared concepts among categories and the model channels. This result substantiates the validity of FS. Besides, extensive experiments on three benchmarks, along with three FSDG variants, show that FSDGs outperform their counterparts in terms of FGDG performance.
@@
The remainder of this article is organized as follows. Section II reviews related works. Problem formulation and the proposed FS method are presented in Section III. In Section IV, the experimental analysis of the proposed method is presented. The explainability analysis is explored in Section V. Finally, Section VI concludes this article with a brief discussion on limitations and future work.
@@
Figure 2: Illustration of the model exemplified in a three-granularity hierarchy (i.e., $G=3$). The FS module is highlighted in the purple box in which the solid arrows depict the operation dimensions. Given five input images, FSDG outputs multi-granular results, as indicated in the right section. $\oplus$ represents the disentanglement operator. The box showing $\mathcal{L}_{lf}$ in the right region illustrates the alignment operation for coarse-fine predicted distributions and shows $\varepsilon$ during the training process. Conf. is an abbreviation of confounding and Granu. Trans. means the Granularity Transition Layer. The coarse branches and the FS optimization module will be excluded during model inference.
@@ II-A Domain Generalization
Robust domain-invariant representations emphasize computing the consistencies across multi-source domains [5]. Rosenfeld et al, [3] acquired domain-invariant features with the help of causal analysis. Literature [29] computed a kind of rationale invariance to enhance DG performance. Data augmentation methods can be used to generate multiple domain data for single-source DG problems [30]. [31] employed adversarial training to generate images with unseen styles. Zhou et al. [23] mixed the statistics of two images in the image or feature level to generate new styles. Wang et al. [25] improved DG capability by implicitly aligning the gradient directions between the empirical risk and the perturbed loss. Zhang et al. [1] introduced a training strategy called Lookahead, which involves weight interpolation, to explore flat minima for DG issues. In [32], two separate encoders were learned in an adversarial way to capture identity and domain information respectively. Low-rank decomposition on weight matrices is applied in [33] to identify features that are more generalizable.
@@ II-B Fine-Grained Visual Categorization
Fine-grained visual categorization (FGVC) strives to discriminate subtle distinctions between subordinate categories within the same root category [6]. It presents a formidable challenge due to the inherent dilemma of minimal inter-class variation coupled with significant intra-class variation. Researchers endeavoring to handle this problem can be classified into three main paradigms: feature enhancement-based [34], localization-based [35] and external knowledge-based methods [36]. These methods mainly focus on enhancing the fine discriminative characteristics. For example, FET-FGVC [34] proposed a feature-enhanced Transformer, cooperating with Graph Neural Networks, to improve the performance of FGVC. Zheng et al, [35] firstly localized key object parts within images and then found discriminative clues on object parts. Sun et al, [36] incorporated object structure knowledge into the transformer to upgrade discriminative representations. As for fine-grained generalization issues, PAN [7] harnessed multi-granularity knowledge aligning the final multi-granular predictions to improve fine-grained DA performance. WDAN [37] utilized a weighted discriminative adversarial network to release the fine-grained DA issue. However, these methods neglect feature invariance between categories, and the single-source FGDG problem has not been explored.
@@ III Method
In this section, we introduce our problem formulation. Meanwhile, we introduce the details of our proposed FSDG approach. We first illustrate how our method disentangles feature space into three parts. Then an aligning strategy is adopted to achieve semantic functionalization.
@@ III-A Problem Statement
Fine-Grained Domain Generalization: In the FGDG problem, we train a model $f$ on $M$ source domains $\mathcal{D_{S}}=\left\{\mathcal{S}^{s}=\left\{\left(x^{s},y_{f}^{s},\left.y_{g}^{s}\right|_{g=1}^{G-1}\right)\right\}\right\}_{s=1}^{M}$ and test it on a target domain $\mathcal{D_{T}}=\left\{\mathcal{T}\right\}$. The label space of $\mathcal{D_{S}}$ is constructed in a multi-granular format with a total $G$-layer hierarchy (i.e., $\mathcal{Y}=\left\{(y_{f},\left.y_{g}\right|_{g=1}^{G-1})\right\}$. Traditional DG can be seen as a particular kind of FGDG whose label space has only the fine-grained labels $y_{f}$ without the coarse-grained labels $\left.y_{g}\right|_{g=1}^{G-1}$, where $g$ is a notation of granularity level and subscript $f$ represents the fine-grained content (i.e., $g=0$). The objective of FGDG is to minimize the fine-grained prediction error on the unseen target domain $\mathcal{D_{T}}$. In this paper, we enhance FGDG performance in a single-source scenario (i.e., $M=1$), taking into account a more realistic situation in which multi-source fine-grained structured data is more difficult to fabricate. Additionally, single-source FGDG can demonstrate the efficacy of our method more effectively.
Accordingly, we construct a baseline FGDG model as shown in Fig. 2. Macroscopically, feature extractors $E$ extract semantic features, followed by granularity transition layers $T$ to filter the corresponding information. Finally, the Global Average Pooling (GAP) and Fully Connected (FC) classifier $C$ predict the resulting distributions. Integrating disentanglement and alignment, the FS optimization module is introduced to structuralize the features systematically.
@@ III-B Disentanglement and Decorrelation of Three Partitions
In this paper, our objective is to enhance FGDG performance. In order to fully exploit the multi-granularity information, we configure two backbones to extract coarse and fine features separately, as shown in Fig. 2. The dedicated fine features are utilized to acquire the finest-grained classification. After two backbones, $G$ transition layers parallelly decompose the coarse and fine backbone features into features unique to each granularity, resulting in $G$ branches working in a multi-task classification manner.
Specifically, given a batch of source domain images $x\in\mathbb{R}^{B\times 3\times W\times H}$, two feature extractors $E_{c}$ and $E_{f}$ extract coarse and fine features $\mathcal{F}_{c}$ and $\mathcal{F}_{f}$, respectively. Then $\left.T_{g}\right|_{g=0}^{G-1}$ take them as input to decompose the granular features $\left.\mathcal{F}_{g}\right|_{g=0}^{G-1}$ for every granularity. For branch $g$, $\mathcal{F}_{g}$ consists of a set of features $\left\{{f}_{g}^{1},{f}_{g}^{2},\cdots,{f}_{g}^{d}\right\}$, where $d$ is the channel number of $T_{g}$. Finally, $\left.C_{g}\right|_{g=0}^{G-1}$ output the predicted distributions belonging to every granularity. It is imperative to elucidate that the auxiliary coarse branches and the entire FS optimization module will be excluded during inference, signifying that the ultimate deployed model comprises exclusively the fine-grained branch. Moreover, the feasibility of training with a solitary backbone is also evident.
Features ${f}_{g}\in\mathbb{R}^{B\times 1\times w\times h}$ of CNN or tokens of Transformer represent distinctive semantic concepts of the input images  [18, 19]. Taking CNN as an example, we obtain $d$ semantic features $\mathcal{F}_{g}$ after $T_{g}$. We disentangle them into three segments:
@@
In this paper, we disentangle features at the conceptual level according to their channel indices as the interpretability analysis shows that channels can express various semantic concepts [18]. This procedure is also characterized by its conciseness, practicality, and computational efficiency. It allows us to explicitly identify the index positions of the functionality of the three parts, enhancing the transparency of the model.
To enhance the degree of feature disentanglement from a semantic perspective, we argue that the semantic conceptual features described in the three segments should be as independent and orthogonal as possible. We first compute the prototypes of the disentangled segments:
@@
where $SUM(X)$ is the summation operator among all elements of tensor $X$. Cosine similarity is utilized as the metric because it can directly quantify orthogonality. Other metrics, such as Hilbert-Schmidt independence criterion (HSIC) [39] and Euclidean distance [40], are also tested, and the results and analyses are presented in Section IV.

### v2

@@ Abstract
Fine-grained domain generalization (FGDG) is a more challenging task than traditional DG tasks due to its small inter-class variations and relatively large intra-class disparities. When domain distribution changes, the vulnerability of subtle features leads to a severe deterioration in model performance. Nevertheless, humans inherently demonstrate the capacity for generalizing to out-of-distribution data, leveraging structured multi-granularity knowledge that emerges from discerning the commonality and specificity within categories. Likewise, we propose a Feature Structuralized Domain Generalization (FSDG) model, wherein features experience structuralization into common, specific, and confounding segments, harmoniously aligned with their relevant semantic concepts, to elevate performance in FGDG. Specifically, feature structuralization (FS) is accomplished through joint optimization of five constraints: a decorrelation function applied to disentangled segments, three constraints ensuring common feature consistency and specific feature distinctiveness, and a prediction calibration term. By imposing these stipulations, FSDG is prompted to disentangle and align features based on multi-granularity knowledge, facilitating robust subtle distinctions among categories. Extensive experimentation on three benchmarks consistently validates the superiority of FSDG over state-of-the-art counterparts, with an average improvement of 6.2% in FGDG performance. Beyond that, the explainability analysis on explicit concept matching intensity between the shared concepts among categories and the model channels, along with experiments on various mainstream model architectures, substantiates the validity of FS.
@@ I Introduction
In FGDG tasks, fine-grained categories exhibit multi-granularity structures built upon the commonalities and specificities between categories. This aspect is overlooked by previous DG methods. In contrast, humans exhibit a higher level of generalizability by leveraging these structured semantics for learning and recognizing. According to family resemblances [9], feature integration theory [10], and lexical structure [11] in cognitive psychology, when learning to recognize objects, people focus primarily on the salient targets and inductively disentangle the commonalities and specificities among them. Commonalities represent basic visual concepts, whereas specificities correspond to individual conceptual extensions. Both serve as the foundational elements for object classification and vary across distinct granularity levels [12, 13, 14]. For instance, as Fig. 1 shows, four animals can be reclassified into categories of Equidae and Bovidae based on their commonalities and specificities from a higher $c_{1}$ granularity level, respectively. At $c_{1}$ level, zebra and horse share certain characteristics, such as long necks, while manifesting specificities at fine-grained level $f_{0}$. It can be observed that the common and specific features, which constitute the intrinsic characteristics of objects, can be facilitated by incorporating multi-granularity knowledge. Understanding the commonalities and specificities between categories enhances the generalizability, as both focus more on the intrinsic features of the objects themselves.
@@
However, a key challenge arises: how to explicitly embed these structured commonalities and specificities into DL models. In particular, it is unclear to determine what semantic basis can be used for constructing clusters of commonalities and specificities, and it is challenging to constrain them within their respective semantic clusters, as literature has discussed the unattainability of learning decoupling semantics in the absence of sufficient additional knowledge constraints [15].
@@
To address this challenge, we propose a Feature Structuralization (FS) framework employing multi-granularity knowledge as an additional constraint for semantic disentanglement and alignment. Researches in the explainability field show that features or tokens behave as distinct semantics [16, 17, 18, 19], which implies that commonalities and specificities can be disentangled from the learned features. Acknowledging that recognition tasks predominantly seek to discern target objects in an image that encompasses target semantics and confounding, we disentangle the total semantical features learned from an image into three components, namely commonality, specificity, and confounding. Furthermore, the representations of commonalities and specificities are constrained within their respective semantic clusters with the aid of a multi-granularity knowledge structure. The collaboration between both facilitates the network in learning category invariance while preserving learning discrimination. Consequently, FS optimizes the fine-grained invariant representations as both focus more on the attributes of the objects themselves rather than on the other spurious correlations.
@@
To formulate FS, we devise a pipeline consisting of two essential steps: disentanglement and alignment. We disentangle the three parts according to their channel indices. A decorrelation optimization function is employed to minimize the mutual information between them. Three similarity-based losses are then proposed for the alignment of common and specific parts with the aid of multi-granularity knowledge. Two of these are leveraged to pull commonalities closer across diverse granularities, while the third serves to create distance among the specific components. Furthermore, a Feature Structuralized Domain Generalization model (FSDG) is constructed by incorporating a prediction calibration method. Three model variations are expanded in conjunction with Convolutional Neural Network (CNN), Transformer, and Multi-Layer Perceptron (MLP).
In summary, the contributions of this work are as follows:
@@
(1) We propose an FS framework to address a more difficult FGDG challenge, drawing inspiration from cognitive psychology. This novel approach integrates data and multi-granularity knowledge to structure the learned features into common, specific, and confounding parts according to their channel indices, which is in line with the analysis of explainability. FS endows the model with the ability to explicitly learn commonalities and specificities between categories.
@@
(2) We construct a feature-structuralized model FSDG by formulating FS into a disentanglement and alignment pipeline. A disentanglement enhancement function is derived to reduce the correlation between the three segments. Three constraints are then designed to achieve semantic alignment for the common and specific components. To our knowledge, FSDG is a pioneer tailored to address FGDG issues.
@@
(3) A CRP method is refined to substantiate the validity of FS. It statistically computes the most relevant concepts for every category, showing that FS significantly increases the explicit concept matching intensity between the shared concepts among categories and the model channels. Besides, extensive experiments on three benchmarks, along with three FSDG variants, show that FSDGs outperform their counterparts in terms of FGDG performance.
@@
The remainder of this article is organized as follows. Section II reviews related works. Problem formulation and the proposed FS framework are presented in Section III. In Section IV, the experimental analysis of the proposed approach is presented. The explainability analysis is explored in Section V. Finally, Section VI concludes this article with a brief discussion on limitations and future work.
@@
Figure 2: Illustration of the model exemplified in a three-granularity hierarchy (i.e., $G=3$). The FS module is highlighted in the middle box with solid arrows depicting the operation dimensions. $\oplus$ represents the disentanglement operator. Given five input images, FSDG outputs multi-granular results, as indicated in the right section. The box showing $\mathcal{L}_{lf}$ illustrates the alignment operation for coarse-fine predicted distributions and shows $\varepsilon$ during the training process. Conf. is an abbreviation of confounding and Granu. Trans. means the Granularity Transition Layer. The coarse branches and the FS optimization module are excluded during model inference.
@@ II-A Domain Generalization
Robust domain-invariant representations emphasize computing the consistencies across multi-source domains [5]. Rosenfeld et al. [3] acquired domain-invariant features with the help of causal analysis. Literature [29] computed a kind of rationale invariance to enhance DG performance. Data augmentation methods can be used to generate multiple domain data for single-source DG problems [30]. [31] employed adversarial training to generate images with unseen styles. Zhou et al. [23] mixed the statistics of two images in the image or feature level to generate new styles. Wang et al. [25] improved DG capability by implicitly aligning the gradient directions between the empirical risk and the perturbed loss. Zhang et al. [1] introduced a training strategy called Lookahead, which involves weight interpolation, to explore flat minima for DG issues. In [32], two separate encoders were learned in an adversarial way to capture identity and domain information, respectively. Low-rank decomposition on weight matrices is applied in [33] to identify features that are more generalizable.
@@ II-B Fine-Grained Visual Categorization
Fine-grained visual categorization (FGVC) strives to discriminate subtle distinctions between subordinate categories within the same root category [6]. It presents a formidable challenge due to the inherent dilemma of minimal inter-class variation coupled with significant intra-class variation. Researchers endeavoring to handle this problem can be classified into three main paradigms: feature enhancement-based [34], localization-based [35], and external knowledge-based methods [36]. These methods mainly focus on enhancing the fine discriminative characteristics. For example, FET-FGVC [34] proposed a feature-enhanced Transformer, cooperating with Graph Neural Networks, to improve the performance of FGVC. Zheng et al. [35] firstly localized key object parts within images and then found discriminative clues on object parts. Literature [36] incorporated object structure knowledge into the transformer to upgrade discriminative representations. As for fine-grained generalization issues, PAN [7] harnessed multi-granularity knowledge, aligning the final multi-granular predictions to improve fine-grained DA performance. WDAN [37] utilized a weighted discriminative adversarial network to release the fine-grained DA issue. However, these methods neglect feature invariance between categories, and the single-source FGDG problem has not been explored.
@@ III Method
In this section, we introduce our problem formulation and the details of our proposed FS framework. We illustrate how our method disentangles feature space into three parts. An aligning strategy is then adopted to achieve semantic functionalization.
@@ III-A Problem Statement
Fine-Grained Domain Generalization: In the FGDG problem, a model $f$ is trained on $M$ source domains $\mathcal{D_{S}}=\left\{\mathcal{S}^{s}=\left\{\left(x^{s},y_{f}^{s},\left.y_{g}^{s}\right|_{g=1}^{G-1}\right)\right\}\right\}_{s=1}^{M}$ and tested on a target domain $\mathcal{D_{T}}=\left\{\mathcal{T}\right\}$. The label space of $\mathcal{D_{S}}$ is constructed in a multi-granular format with a total $G$-layer hierarchy (i.e., $\mathcal{Y}=\left\{(y_{f},\left.y_{g}\right|_{g=1}^{G-1})\right\}$). Traditional DG can be seen as a particular kind of FGDG whose label space has only the fine-grained labels $y_{f}$ without the coarse-grained labels $\left.y_{g}\right|_{g=1}^{G-1}$, where $g$ is a notation of granularity level and subscript $f$ represents the fine-grained content (i.e., $g=0$). The objective of FGDG is to minimize the fine-grained prediction error on the unseen target domain $\mathcal{D_{T}}$. In this paper, we enhance FGDG performance in a single-source scenario (i.e., $M=1$), taking into account a more realistic situation in which multi-source fine-grained structured data is more difficult to fabricate. Additionally, single-source FGDG can demonstrate the efficacy of our method more effectively.
Accordingly, we construct a baseline FGDG model as shown in Fig. 2. Macroscopically, feature extractors $E$ extract semantic features, followed by granularity transition layers $T$ to filter the corresponding granularity information. Finally, the Global Average Pooling (GAP) and Fully Connected (FC) classifier $C$ predict the resulting distributions. Integrating disentanglement and alignment, the FS optimization module is introduced to structuralize the features systematically.
@@ III-B Disentanglement and Decorrelation of Three Partitions
In this paper, our objective is to enhance FGDG performance utilizing multi-granularity knowledge. In order to fully exploit this knowledge, we configure two backbones to extract coarse and fine features separately, as shown in Fig. 2. $G$ transition layers parallelly decompose the coarse and fine backbone features into features unique to each granularity, resulting in $G$ branches working in a multi-task manner. The fine branch is dedicated to identifying the fine-grained categories, while the features of coarse branches assist the fine-grained features in capturing commonalities and specificities.
Specifically, given a batch of source domain images $x\in\mathbb{R}^{B\times 3\times W\times H}$, two feature extractors $E_{c}$ and $E_{f}$ extract coarse and fine features $\mathcal{F}_{c}$ and $\mathcal{F}_{f}$, respectively. Then $\left.T_{g}\right|_{g=0}^{G-1}$ take them as input to decompose the granular features $\left.\mathcal{F}_{g}\right|_{g=0}^{G-1}$ for every granularity. For branch $g$, $\mathcal{F}_{g}$ consists of a set of features $\left\{{f}_{g}^{1},{f}_{g}^{2},\cdots,{f}_{g}^{d}\right\}$, where $d$ is the channel number of $T_{g}$. Finally, $\left.C_{g}\right|_{g=0}^{G-1}$ output the predicted logits belonging to every granularity. It is imperative to elucidate that the auxiliary coarse branches and the entire FS optimization module will be excluded during inference, signifying that the ultimate deployed model comprises exclusively the fine-grained branch. Moreover, the feasibility of training with a solitary backbone is also evident.
Features ${f}_{g}\in\mathbb{R}^{B\times 1\times w\times h}$ of CNN or tokens of Transformer represent distinctive semantics of the input images  [18, 19]. Taking CNN as an example, we obtain $d$ semantic features $\mathcal{F}_{g}$ after $T_{g}$. We disentangle them into three segments:
@@
In this paper, we disentangle features at the conceptual level according to their channel indices, as the interpretability analysis shows that channels can express various concepts [18]. This procedure is also characterized by its conciseness, practicality, and computational efficiency. It allows us to explicitly identify the index positions of the functionality of the three parts, fortifying the transparency of the model.
To enhance the degree of disentanglement from a semantic perspective, we argue that the conceptual features described in the three segments should be as independent and orthogonal as possible. We first compute the prototypes of the segments:
@@
where $SUM(X)$ is the summation operator among all elements of tensor $X$, and $\mathbf{I}$ is a identity matrix. Cosine similarity is utilized as the metric because it can directly quantify orthogonality. Other metrics, such as Hilbert-Schmidt independence criterion (HSIC) [39] and Euclidean distance [40], are also tested, and results are presented in Section IV.